OpenStudy (anonymous):

integrate (t^-2)(e^t)dt

OpenStudy (anonymous):

use integration by parts.

OpenStudy (anonymous):

how exactly?

OpenStudy (anonymous):

int (u *dv) = uv- int (v * du) in this case, you let u=t^-2 and v= e^t

OpenStudy (anonymous):

I tried that. doesn't work because du will become -2t^-3, which doesn't help.

OpenStudy (anonymous):

so try dv=t^-2 and u=e^t. u may need to do integration by parts twice.

OpenStudy (anonymous):

i tried that too. eventually the integral becomes lnt*e^tdt. How do you integrate that?

OpenStudy (anonymous):

i don't think you can integrate this and get a transdencental function

OpenStudy (anonymous):

thanks everyone.